ABCD is a square and P, Q are the midpoints of BC, CD respectively. If AP = a and AQ = b, find in terms of a and b, the directed line segments (i) AB, (ii) AD, (iii) BD and (iv) AC.


1
Expert's answer
2021-01-25T09:53:30-0500


To find AB and AD we have to make some equations. First let us sign BP = x, AD = 2x, QD = y, AB = 2y. From the triangle ABP and from triangle AQD we can write two equations using the Pythagoras Theorem:  

"4y^2+x^2=a^2"

"4x^2+y^2=b^2"


"x^2=a^2\u22124y^2"

"4(a^2\u22124y^2)+y^2=b^2"


"4a^2\u221216y^2+y^2=b^2"


"4a^2\u221215y^2=b^2"


"y^2=(4a^2\u2212b^2)\/15"


"y=\sqrt{(4a^2−b^2)/15} ; 2y=2\sqrt{(4a^2−b^2)/15}"


"x^2=a^2\u22124(4a^2\u2212b^2)\/15"


"x^2=(15a^2\u221216a^2+4b^2)\/15"


"x^2=(4b^2\u2212a^2)\/15"


"x=\sqrt{(4b^2−a^2)/15} ; 2x=2\sqrt{(4b^2−a^2)/15}"


So "AB=2y=2\sqrt{(4a^2−b^2)/15}"


"AD=2x=2\sqrt{(4b^2−a^2)/15}"


From the triangle ABD we can find BD using the Pythagoras Theorem: 


"BD=AC=2\sqrt{(4a^2−b^2)/15+(4b^2−a^2)/15}="


"=2\sqrt{(4a^2−b^2+4b^2−a^2)/15}=2\sqrt{(3a^2+3b^2)/15}=2\sqrt{(a^2+b^2)/5}"